OpenStudy (anonymous):

Which sums are equal to 1 2/6

OpenStudy (study312):

multiple choice?

rebeccaxhawaii (rebeccaxhawaii):

^^^

OpenStudy (anonymous):

choices?

OpenStudy (study312):

he is offline

OpenStudy (anonymous):

he? you mean she

OpenStudy (anonymous):

6+/6+2/6=

OpenStudy (anonymous):

the other one was a fool

OpenStudy (study312):

1 • 6 + 2 = 8/6